[figure: 9.7 cm]

SPECIFIC HEAT STUDIES OF S=1 HEISENBERG PLANAR ANTIFERROMAGNETIC CHAIN Ni(C_{11}H_{10}N_{2}O)_{2}Ni(CN)_{4}

E. Čižmár, J. Černák, M. Orendáč and A. Feher

Faculty of Science, P.J. Šafárik University, Park Angelinum 9, 04154 Košice, Slovakia

In this work we discuss the results of experimental studies of specific heat of chain compound Ni(C_{11}H_{10}N_{2}O)_{2}Ni(CN)_{4}, which were performed in the temperature range 100mK-10K. We have observed a Schottky type anomaly at T = 2.2K in temperature dependence of specific heat and the compound was identified as S=1 Heisenberg antiferromagnetic chain with subcritical exchange coupling D/k_{B} = 5K, D/|J| = 11 and strong in-plane anisotropy E/k_{B} = 2.75K. It has been found that in-plane anisotropy affects on the region of validity of a diluted excitons model [1] and strong coupling model [2] for S=1 Heisenberg chains with strong planar anisotropy. The analysis suggests that in-plane anisotropy should be considered in any attempt to find a compound potentially located in the boundary of Haldane and ”large-D” phases.

[1] N. Papanicolaou, P. Spathis, J. Phys. Condens. Matter 2 (1990), 6575.

[2] N. Papanicolaou, P. Spathis, Phys. Rev. B 52 (1995), 16001.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Erik Čižmár

Address for correspondence : Department of Experimental Physics Faculty of Science, P.J. Šafárik University Park Angelinum 9 04154 Košice Slovakia

Email address : cizmar@kosice.upjs.sk